Log in
wiki.ucar.eduwiki.ucar.edu
  • Spaces
  • Collections
  • Hit enter to search
  • Help
    • Online Help
    • Keyboard Shortcuts
    • Feed Builder
    • What’s new
    • About Confluence
  • Log in

METCRAXII Field Project - ISFS
  • Pages
  • Blog

Space shortcuts

  • University of Utah Metcrax2 page
  • EOL METCRAXII web page
  • ISS Logbook
  • Data Tables and Plots

Blog

  • 2014
  • 2013
    • November
    • October
      • Teardown status soils boom angles
      • the end has come
      • teardown starting
      • Daily status 29 October
      • Daily status 28 October
      • visit to flr
      • odd w.3m.near
      • post mortem issues
      • Daily status 27 October
      • Generator
      • flr locations
      • flr servicing
      • cleaning far
      • cleaning near
      • Daily status 26 October
      • USBdisk at NEAR went down
      • Daily Status 25 October
      • TRH.40m.rim restarted
      • kh2o flr glitch
      • Qsoil at base
      • Daily Status 24 October
      • playing with mote power
      • Qsoil offset adjustments
      • SPN1 thoughts
      • P.nne.flr issues
      • Krypton at floor cleaned
      • P.sw.flr back up
      • Daily Status 23 October
      • P.sw down
      • Generator Change
      • Soil moisture sensing
      • Daily status 22 October
      • flr visit
      • Radiometers cleaned
      • first gravimetric samples
      • Daily status 21 October
      • near SPN1 mote died again
      • Vpile.off range expanded
      • trailer WiFi settings
      • Daily status 20 October
      • NNE servicing
      • fixed spn1.near mote
      • flr cleanings
      • Generator fueling
      • Daily status 19 October
      • new metcraxii config
      • NEAR TRH work
      • Near mote work
      • Mote Near - Far xbee reset rates and 'hb'
      • glitch on kh2o far
      • Daily status, October 18
      • near rads cleaned
      • external network outage
      • P.nne down yesterday
      • Daily status, October 17
      • flr network outages
      • Changed power monitor mote ID - Near
      • flr motes set to serial mode
      • Mote Xbee status message rate disabled
      • cover_redo with new sonic tilt angles
      • Daily status, October 16
      • flr Mote check - restarts
      • Service visit to NEAR
      • Service visit to FAR
      • FLR mote heartbeats every 1 hour
      • Daily status, October 15
      • Pulled FAR Rlw.out
      • Replaced 20m RIM TRH
      • Processing notes
      • Notes for use of R at Metcraxii
      • FLR servicing
      • Daily status, October 14
      • Generator refilled at ISFS base trailer
      • fix inadyn issue
      • Added periodic send of heart-beat command to base motes
      • Daily status, October 13
      • Service visits to NEAR
      • Service visit to FAR
      • Service visit to RIM
      • Guy wire tensions
      • Swapped barometers in crater, October 11
      • Power monitor motes
      • Reseated Rsw.in rad shield at FAR
      • Service visit to NEAR
      • Service visit to FAR
      • Voltage level of crater stations
      • Service visit to crater floor Oct 11
      • Daily status, October 10
      • Mote radio reset, gps sync timing adjusted
      • Generator refueled
      • 20 and 40m TRHs on RIM down
      • SPN1 high
      • P glitches
      • NEAR and RIM wind direction adjustment
      • Ifan values 0 at 25m.near, 20m.rim
      • new UPS, reworked trailer network
      • spectra of kryptons
      • Daily status October 9
      • Update to Solar panel loan to German group
      • Service visit to FAR
      • Service visit to NEAR
      • Color wind plot
      • (Final) sonic tilts
      • Daily status, October 8
      • soil comments
      • Generator service
      • mote status
      • Daily status October 6-7
      • Loaning backup solar panels to Norbert
      • Changes to repeater, added to rim tower battery pack
      • Soil Install Thoughts
      • Guy Wire Tensions
      • Network outage in ISFS trailer
      • Krypton at FLR
      • Service visit to NEAR
      • 10m wind direction at FAR and FLR
      • Service visit to RIM
      • Service visit to FAR
      • Service visit to NEAR
      • mote repair
      • 2D sonic azimuth at base
      • Daily status, Oct 5
      • FLR krypton hygrometer
      • FLR radiation mote
      • FLR sonic orientation
      • SSW2 and SSW4 sonics
      • Trailer power brownout caused AP24 to go down
      • RIM data system down
      • Status, October 3
      • Replaced power cable on radiation mote at FAR
      • Flash local storage at near
      • WIFI status, enable internal antenna at base
      • FAR maintenance
      • NEAR maintenance
      • RIM Tower Sonic Heights
    • September
    • August
    • June
Browse pages
  • View Page
    • Attachments (2)
    • Page History
    • Page Information
    • View Source
    • Export to PDF
    • Hide Inline Comments
  1. Dashboard
  2. METCRAXII Field Project - ISFS
  3. 2013
  4. October
  5. 09
  6. spectra of kryptons
  7. Attachments
  8. krypton_spectra.pdf

krypton_spectra.pdf

Overview
Content Tools

Powered by a free Atlassian Confluence Community License granted to University Corporation for Atmospheric Research. Evaluate Confluence today.

  • Powered by Atlassian Confluence 9.2.13
  • Printed by Atlassian Confluence 9.2.13
  • Report a bug
  • Atlassian News
Atlassian
{"serverDuration": 99, "requestCorrelationId": "4556d6e9d158f87d"}